# references.bib

### v1

@@ 5.1.1. Direct Evaluations
As a way of introducing Intel CAT to the scientific community, herdrich_cache_2016 (herdrich_cache_2016) present a feature showcase in their (herdrich_cache_2016) paper. The publication addresses the specifics of the implementation and highlights the importance of the feature by suggesting use cases for the new technology. To back up their claims, the effect on the performance of workloads constrained by LLC allocations is presented and discussed.
@@
In their recent study, sohal_closer_2022 (sohal_closer_2022) conduct an experimental evaluation of Intel’s CAT and MBA. It is found that the behavior of the features differs between chip generations. Therefore, they suggest using their evaluation approach for coming architectural versions in order to verify the behavior of such versions.
@@
zhang_libra_2021 (zhang_libra_2021) conducts a smaller study that highlights the shortcomings of Intel MBA. This evaluation is used as motivation when introducing their own software-based approach to memory bandwidth allocation.
@@
Through an experimental analysis, farina_assessing_2022 (farina_assessing_2022) evaluate the effectiveness and behavior of MBA for resource limitation in real-time systems. They conclude that the hardware-based approach to memory bandwidth limiting outperforms similar software-based approaches such as MemGuard (yun_memguard_2013).
@@
In their master thesis, ihre_sherif_evaluation_2021 (ihre_sherif_evaluation_2021) evaluate the performance implications of LLC partitioning on three commonly used workloads: bzip2, redis, and Graph500. They find that tuned partitioning may increase the performance of the aforementioned workloads by up to $5.8\%$, $8.6\%$, and $12.0\%$, respectively, compared to an unmanaged system when co-located with a noisy neighbor.
@@
schramm_impact_2019 (schramm_impact_2019) explore the effectiveness of CAT as a method of improving the performance of virtual network functions (VNFs). They find that sufficient LLC allocation may increase the throughput of a VNF by up to $33\%$ during heavy load scenarios.
@@
To better understand the implications of cache partitioning on workloads with different characteristics, qin_when_2018 (qin_when_2018), kim_simple_2017 (kim_simple_2017), and veitch_performance_2017 (veitch_performance_2017) conduct empirical studies where the behavior of a range of workloads under different cache partitioning schemes is examined. They find that the tuning of cache allocation schemes is generally beneficial to ensuring predictable performance of workloads.
@@
While not strictly presenting a feature evaluation, castiglione_optimal_2018 discuss the potential usefulness of cache allocation in preventing and mitigating cache side-channel attacks (castiglione_optimal_2018). They observe the behavior enabling the side-channel attack vector and argue for the applicability of cache partitioning as a possible solution to the problem. However, determining if cache allocation upholds the promises is left as future work.
@@ 5.1.2. Indirect Evaluations
With the rise of serverless computing and function-as-a-service (FaaS) platforms, an increasing number of practitioners are looking to the promises of ease-of-use and pay-per-use of the paradigm. qiu_is_2021 (qiu_is_2021) recognize the need to evaluate the feasibility of FaaS as a model to host LC workloads. In their evaluation, they include LLC and memory bandwidth partitioning as potential ways of ensuring consistent performance. Their findings conclude that FaaS is not yet ready for LC workloads and that the aforementioned partitioning techniques would have to support orders of magnitude more CLOSs before QoS of FaaS functions can be sufficiently enforced.
On platforms with integrated graphics, LLC is often shared between the CPU and GPU. This becomes an attack vector for denial-of-service (DoS) attacks. bechtel_denial--service_2022 (bechtel_denial--service_2022) evaluates if LLC partitioning using CAT provides sufficient protection to mitigate such attacks. They conclude that CAT may indeed provide spatial isolation, but that there are still cases where temporal isolation can not be achieved. This in turn leads to sophisticated attacks still being viable.
To better understand the mechanics of Intel’s Data Direct I/O (DDIO) and its correlation to LLC cache ways, farshin_reexamining_2020 (farshin_reexamining_2020) conducts an empirical study to gain a better understanding of the inner workings of the technology. With the knowledge gained from this study, they discuss its effectiveness and shortcomings. General advice regarding how DDIO should be utilized depending on the use case is also presented.
shang_coplace_2021 (shang_coplace_2021) utilize LLC partitioning to demonstrate that caching inefficiencies exist in the way hosts map guest VM memory pages to physical memory. Their proposed solution, CoPlace mitigates this problem and reduces cache conflict misses that exist even when a guest is given exclusive access to some subset of LLC.
@@ 5.2. Black-Box Approaches
The ability to tune a system without knowledge of the applications it is hosting can be useful. However, doing so may be challenging considering that different applications are sensitive to different bottlenecks (cortez_resource_2017). Even so, it may be required in some cases due to privacy concerns. Such is the case with public cloud providers that are unable to observe the workloads running in customers’ virtual machines (shahrad_provisioning_2021). Consequently, tuning must be done by observation of system metrics only. Thus, tuning becomes workload agnostic, as all processes are observed using the same system and operating system metrics.
Proposals leveraging the surveyed platform features that treat workloads as “black boxes” are presented in Table 4. We split black box approaches into a number of subclasses: Proactive approaches, reactive approaches, combined approaches, workload clustering approaches, and feature abstractions.
@@
(xiao_cppf_2019) Generic Performance Prefetch-Aware Heuristic LLC Partitioning
(kim_application_2019) Generic Perf. Prediction ML Performance Prediction Model
(aupy_co-scheduling_2019) HPC Perf. Prediction Mathematical Modelling of Execution Times
(durner_towards_2019) VNFs Utilization Static Heuristic Optimization Algorithm
(funaro_ginseng_2016) Cloud Performance Market-Driven Auction-Based Scheme
(shahrad_provisioning_2021) Cloud Performance Cache-Aware VM Scheduler
(liu_catalyst_2016) Security Security LLC Side-channel Attack Protection Scheme
(dessouky_chunked-cache_2021) Security Perf., Security LLC Side-channel Attack Protection Scheme
(zeng_playing_2021) Generic Performance Greedy Static LLC Allocation Scheme
(chung_enforcing_2019) Generic Performance LLC-Aware Memory Bandwidth Management
(thomas_dark_2018) VNFs Performance Memory Hierarchy Management Framework
(sprabery_novel_2017) Security Security LLC Side-Channel Attack Protection Scheme
@@
(navarro-torres_balancer_2023) Generic Perf., Fairness Heuristic LLC and MB Management Scheme
(park_hypart_2018) Generic Performance Explorative MB Partitioning Algorithm
(park_copart_2019) Generic Perf., Fairness Explorative LLC and MB Partitioning
(xu_dcat_2018) Generic Performance Heuristic Program Phase Detection
(pons_cache-poll_2022) Generic Perf., Fairness Heuristic LLC Partitioning Scheme
(farina_enabling_2023) RT Performance Heuristic Bandwidth Guarantees
(roy_satori_2021) Generic Perf., Fairness Bayesian Optimization Based Algorithm
(chen_alita_2020) Cloud Fairness Heuristic VM Resource Fairness Algorithm
(xiang_dcaps_2018) Generic Performance Simulated Annealing Optimization
(yuan_dont_2021) Generic Performance Heuristic, DDIO Aware LLC Management
(yi_mt2_2022) Generic Performance Heuristic MB Management Algorithm
(yao_leveraging_2019) Security Security Cache Timing Channel Attack Mitigation
@@
(li_pfa_2022) Generic Perf., Fairness Heuristic LLC Management Algorithm
(garcia-garcia_lfoc_2019) Generic Fairness Heuristic Cache-Clustering Algorithm
(saez_lfoc_2022) Generic Fairness Heuristic Cache-Clustering Algorithm
(pons_improving_2018) Generic Perf., Fairness Heuristic LLC Partitioning Scheme
(pons_phase-aware_2020) Generic Perf., Fairness Heuristic LLC Partitioning Scheme
(chen_drlpart_2021) Generic Performance Deep Reinforcement Learning Model
(chatterjee_com-cas_2023) Generic Performance Compiler Guided LLC Management
(zhang_zeus_2021) Cloud Utilization Kubernetes Resource Management Scheme
(zhu_kelp_2019) ML Performance Data Management for ML Systems
@@
(tang_themis_2023) Generic Perf., Fairness K-Means Classification of Guest VMs
(gifford_dna_2021) Generic Performance Workload Clustering Technique
(xiang_emba_2019) Generic Performance Heuristic Hierarchical Clustering Algorithm
(selfa_application_2017) Generic Fairness K-Means Classification of Workloads
(qiu_machine-learning-based_2022) Cloud Performance SVM Classification and BO LLC Partitioning
@@
(xu_vcat_2017) VMs Performance Virtualization of CAT
@@ 5.2.1. Proactive Approaches
Publications featuring approaches classified as proactive attempts to make resource management decisions ahead of time. They may enforce static partitioning schemes or try to predict future resource requirements of workloads.
@@
By profiling workloads as sensitive, or non-sensitive, to hardware prefetching of memory, xiao_cppf_2019 (xiao_cppf_2019) propose a controller that aims to improve LLC management on modern systems equipped with hardware prefetching. They call their approach CPpf, which partitions LLC in accordance with the determined workload classifications. The system allocates a single exclusive cache way to each prefetching sensitive workload, leaving the rest to be shared among non-prefetching sensitive workloads.
@@
With the goal of predicting how a process’ instructions per cycle (IPC) metric changes as a result of a change in its available LLC, kim_application_2019 (kim_application_2019) propose POCAT. POCAT is a trained machine learning model that can predict changes to IPC with an error of 4.7% without affecting current LLC allocations.
@@
To get a better understanding of how runtimes of BE high-performance computing (HPC) workloads are affected by core and LLC allocations, aupy_co-scheduling_2019 (aupy_co-scheduling_2019) create a mathematical model that predicts how job execution times are affected by changes in scarce shared resource availability. Using their model, they are able to more efficiently partition resources among workloads to increase the efficiency of a system.
@@
A simple heuristic algorithm to partition LLC among VNFs is proposed by durner_towards_2019 (durner_towards_2019). Their algorithm distributes available cache ways based on CPU utilization, which results in lowered total system utilization, increasing energy efficiency and QoS. However, only static configurations are considered where virtual network functions (VNFs) and their loads do not change. This makes their algorithm valid for very specific use cases only.
@@
In an attempt to make LLC partitioning more effective, Ginseng (funaro_ginseng_2016) allow consumers (i.e. workloads) to bid on LLC space using Vickrey-Clarke-Groves auctioning. This allocates LLC to the process that would see the largest economic benefit from having said space. As the auctioning and allocation cycle is kept short, processes may acquire more LLC space at critical phases in its operation, while seceding it at times of less sensitivity.
@@
Public cloud providers aim to reduce their total cost of hardware ownership by optimizing its utilization. Making a VM scheduler LLC aware allows the provider to give LLC QoS guarantees to their customers as LLC ways may be exclusively, or partially exclusively, allocated to specific VMs. This is the goal of CacheSlicer, a VM scheduler proposed by shahrad_provisioning_2021 (shahrad_provisioning_2021) that schedules VMs onto nodes based on low-cost, normal, and cache-sensitive VM categories.
@@
In public clouds, where different clients’ VMs share physical hosts, malicious actors may compromise security by extracting sensitive information from a co-located guest by manipulation of the LLC. CATalyst, as proposed by liu_catalyst_2016 (liu_catalyst_2016), is a module that integrates into hypervisors that protect against such LLC side-channel attacks by partitioning the LLC to provide security guarantees.
To improve general performance in trusted execution environments, dessouky_chunked-cache_2021 propose a hybrid cache management scheme that splits applications based on their sensitivity to LLC side-channel attacks. Their approach, titled Chunked-Cache (dessouky_chunked-cache_2021), enforces cache security for sensitive applications by placing them in an isolated domain. One such domain is created and enforced for each sensitive context, while the operating system and other insensitive workloads share a non-sensitive domain which improves overall system performance to similar side-channel protection schemes.
@@
zeng_playing_2021 find that a simple static allocation of LLC is not sufficient to improve system performance when software prefetching of memory is used (zeng_playing_2021). They show that improvements can be made even with a simple greedy partitioning scheme.
@@
An observation that allocating more LLC to an LC application may actually bring worse performance is made by chung_enforcing_2019 (chung_enforcing_2019). This is due to the memory controller being overwhelmed by the requests of the remaining workloads that miss LLC more often due to less remaining usable after partitioning. The paper details their approach to mitigate this issue and the “memory virtual channels” they use to prioritize and order memory requests such that the LC application is guaranteed prioritized access to main memory as well as LLC.
@@
thomas_dark_2018 observe that as networks approach transfer speeds of 100 gigabits per second and beyond, the memory system becomes a bottleneck that can a cause large number of packet drops if incorrectly utilized (thomas_dark_2018). As an attempt to mitigate this issue, they propose CacheBuilder. Cache builder is a system that, among other techniques, utilizes cache allocation to provide applications with limited control of the memory hierarchy such that packet drops may be minimized.
@@
By splitting LLC into two partitions, sprabery_novel_2017 (sprabery_novel_2017) attempts to protect sensitive applications against cache side-channel attacks. Sensitive workloads are assigned to one, while other, insensitive, workloads share the other. The cores assigned to the sensitive partition will upon context switching perform a “state cleansing”. This means that the LLC used by the previous execution context will be flushed before the new context begins execution. This approach protects sensitive data at the cost of performance.

### v2

@@ 5.1.1. Direct Evaluations
As a way of introducing Intel CAT to the scientific community, (herdrich_cache_2016, 41) (herdrich_cache_2016, 41) present a feature showcase in their 2016 paper. The publication addresses the specifics of the implementation and highlights the importance of the feature by suggesting use cases for the new technology. To back up their claims, the effect on the performance of workloads constrained by LLC allocations is presented and discussed.
@@
In their recent study, (sohal_closer_2022, 95) (sohal_closer_2022, 95) conduct an experimental evaluation of Intel’s CAT and MBA. It is found that the behavior of the features differs between chip generations. Therefore, they suggest using their evaluation approach for coming architectural versions in order to verify the behavior of such versions.
@@
(zhang_libra_2021, 122) (zhang_libra_2021, 122) conducts a smaller study that highlights the shortcomings of Intel MBA. This evaluation is used as motivation when introducing their own software-based approach to memory bandwidth allocation.
@@
Through an experimental analysis, (farina_assessing_2022, 27) (farina_assessing_2022, 27) evaluate the effectiveness and behavior of MBA for resource limitation in real-time systems. They conclude that the hardware-based approach to memory bandwidth limiting outperforms similar software-based approaches such as MemGuard (yun_memguard_2013, 118).
@@
In their master thesis, (ihre_sherif_evaluation_2021, 45) (ihre_sherif_evaluation_2021, 45) evaluate the performance implications of LLC partitioning on three commonly used workloads: bzip2, redis, and Graph500. They find that tuned partitioning may increase the performance of the aforementioned workloads by up to $5.8\%$, $8.6\%$, and $12.0\%$, respectively, compared to an unmanaged system when co-located with a noisy neighbor.
@@
(schramm_impact_2019, 91) (schramm_impact_2019, 91) explore the effectiveness of CAT as a method of improving the performance of virtual network functions (VNFs). They find that sufficient LLC allocation may increase the throughput of a VNF by up to $33\%$ during heavy load scenarios.
@@
To better understand the implications of cache partitioning on workloads with different characteristics, (qin_when_2018, 84) (qin_when_2018, 84), (kim_simple_2017, 54) (kim_simple_2017, 54), and (veitch_performance_2017, 103) (veitch_performance_2017, 103) conduct empirical studies where the behavior of a range of workloads under different cache partitioning schemes is examined. They find that the tuning of cache allocation schemes is generally beneficial to ensuring predictable performance of workloads.
@@
While not strictly presenting a feature evaluation, (castiglione_optimal_2018, 30) discuss the potential usefulness of cache allocation in preventing and mitigating cache side-channel attacks (castiglione_optimal_2018, 30). They observe the behavior enabling the side-channel attack vector and argue for the applicability of cache partitioning as a possible solution to the problem. However, determining if cache allocation upholds the promises is left as future work.
@@ 5.1.2. Indirect Evaluations
With the rise of serverless computing and function-as-a-service (FaaS) platforms, an increasing number of practitioners are looking to the promises of ease-of-use and pay-per-use of the paradigm. (qiu_is_2021, 86) (qiu_is_2021, 86) recognize the need to evaluate the feasibility of FaaS as a model to host LC workloads. In their evaluation, they include LLC and memory bandwidth partitioning as potential ways of ensuring consistent performance. Their findings conclude that FaaS is not yet ready for LC workloads and that the aforementioned partitioning techniques would have to support orders of magnitude more CLOSs before QoS of FaaS functions can be sufficiently enforced.
On platforms with integrated graphics, LLC is often shared between the CPU and GPU. This becomes an attack vector for denial-of-service (DoS) attacks. (bechtel_denial–service_2022, 8) (bechtel_denial–service_2022, 8) evaluates if LLC partitioning using CAT provides sufficient protection to mitigate such attacks. They conclude that CAT may indeed provide spatial isolation, but that there are still cases where temporal isolation can not be achieved. This in turn leads to sophisticated attacks still being viable.
To better understand the mechanics of Intel’s Data Direct I/O (DDIO) and its correlation to LLC cache ways, (farshin_reexamining_2020, 29) (farshin_reexamining_2020, 29) conducts an empirical study to gain a better understanding of the inner workings of the technology. With the knowledge gained from this study, they discuss its effectiveness and shortcomings. General advice regarding how DDIO should be utilized depending on the use case is also presented.
(shang_coplace_2021, 94) (shang_coplace_2021, 94) utilize LLC partitioning to demonstrate that caching inefficiencies exist in the way hosts map guest VM memory pages to physical memory. Their proposed solution, CoPlace mitigates this problem and reduces cache conflict misses that exist even when a guest is given exclusive access to some subset of LLC.
@@ 5.2. Black-Box Approaches
The ability to tune a system without knowledge of the applications it is hosting can be useful. However, doing so may be challenging considering that different applications are sensitive to different bottlenecks (cortez_resource_2017, 22). Even so, it may be required in some cases due to privacy concerns. Such is the case with public cloud providers that are unable to observe the workloads running in customers’ virtual machines (shahrad_provisioning_2021, 93). Consequently, tuning must be done by observation of system metrics only. Thus, tuning becomes workload agnostic, as all processes are observed using the same system and operating system metrics. This is illustrated in Figure 9.
Figure 9. Illustration of black-box approaches and their reliance on system and hardware-only metrics. Such metrics may include cache miss rates, memory bandwidth usage, and instructions per cycle numbers, for example. System metrics are generic and exist for all workloads.:
Proposals leveraging the surveyed platform features that treat workloads as “black boxes” are presented in Table 4. Black box approaches are split into subclasses: Proactive approaches, reactive approaches, combined approaches, workload clustering approaches, and feature abstractions.
@@
(xiao_cppf_2019, 110) Generic Performance Prefetch-Aware Heuristic LLC Partitioning
(kim_application_2019, 55) Generic Perf. Prediction ML Performance Prediction Model
(aupy_co-scheduling_2019, 5) HPC Perf. Prediction Mathematical Modelling of Execution Times
(durner_towards_2019, 26) VNFs Utilization Static Heuristic Optimization Algorithm
(funaro_ginseng_2016, 31) Cloud Performance Market-Driven Auction-Based Scheme
(shahrad_provisioning_2021, 93) Cloud Performance Cache-Aware VM Scheduler
(liu_catalyst_2016, 60) Security Security LLC Side-channel Attack Protection Scheme
(zeng_playing_2021, 119) Generic Performance Greedy Static LLC Allocation Scheme
(chung_enforcing_2019, 20) Generic Performance LLC-Aware Memory Bandwidth Management
(thomas_dark_2018, 101) VNFs Performance Memory Hierarchy Management Framework
(sprabery_novel_2017, 96) Security Security LLC Side-Channel Attack Protection Scheme
@@
(navarro-torres_balancer_2023, 67) Generic Perf., Fairness Heuristic LLC and MB Management Scheme
(park_hypart_2018, 75) Generic Performance Explorative MB Partitioning Algorithm
(park_copart_2019, 74) Generic Perf., Fairness Explorative LLC and MB Partitioning
(xu_dcat_2018, 111) Generic Performance Heuristic Program Phase Detection
(pons_cache-poll_2022, 78) Generic Perf., Fairness Heuristic LLC Partitioning Scheme
(farina_enabling_2023, 28) RT Performance Heuristic Bandwidth Guarantees
(roy_satori_2021, 89) Generic Perf., Fairness Bayesian Optimization Based Algorithm
(chen_alita_2020, 15) Cloud Fairness Heuristic VM Resource Fairness Algorithm
(xiang_dcaps_2018, 108) Generic Performance Simulated Annealing Optimization
(yuan_dont_2021, 117) Generic Performance Heuristic, DDIO Aware LLC Management
(yi_mt2_2022, 115) Generic Performance Heuristic MB Management Algorithm
(yao_leveraging_2019, 114) Security Security Cache Timing Channel Attack Mitigation
@@
(li_pfa_2022, 58) Generic Perf., Fairness Heuristic LLC Management Algorithm
(garcia-garcia_lfoc_2019, 33) Generic Fairness Heuristic Cache-Clustering Algorithm
(saez_lfoc_2022, 90) Generic Fairness Heuristic Cache-Clustering Algorithm
(pons_improving_2018, 80) Generic Perf., Fairness Heuristic LLC Partitioning Scheme
(pons_phase-aware_2020, 79) Generic Perf., Fairness Heuristic LLC Partitioning Scheme
(chen_drlpart_2021, 16) Generic Performance Deep Reinforcement Learning Model
(chatterjee_com-cas_2023, 12) Generic Performance Compiler Guided LLC Management
(zhang_zeus_2021, 121) Cloud Utilization Kubernetes Resource Management Scheme
(zhu_kelp_2019, 126) ML Performance Data Management for ML Systems
@@
(tang_themis_2023, 98) Generic Perf., Fairness K-Means Classification of Guest VMs
(gifford_dna_2021, 34) Generic Performance Workload Clustering Technique
(xiang_emba_2019, 109) Generic Performance Heuristic Hierarchical Clustering Algorithm
(selfa_application_2017, 92) Generic Fairness K-Means Classification of Workloads
(qiu_machine-learning-based_2022, 87) Cloud Performance SVM Classification and BO LLC Partitioning
@@
(xu_vcat_2017, 113) VMs Performance Virtualization of CAT
@@
As an example of an important black-box approach, CoPart (park_copart_2019, 74) aims to enhance both performance and fairness of system workloads by classifying them as either sensitive or insensitive to cache and memory bandwidth based on profiling. This classification enables the controller to construct two finite state machines (FSMs) for each application, representing its cache and memory bandwidth characteristics, respectively (see Figure 10(a)). Each FSM contains three states: Supply, where the application can have its allocation reduced without significant performance impact; Demand, where performance would improve if allocated additional resources; And maintain, if performance would degrade if the allocation is reduced, but would only marginally improve if increased. Using these FSMs, CoPart explores the available states to achieve an optimized resource allocation scheme based on current system metrics and by mapping against the Hospitals/Residents (HR) problem (gale_college_1962, 32). The execution flow of CoPart is illustrated in Figure 10(b).
((a)) CoPart resource FSM illustration.:
((b)) CoPart controller execution flow.:
Figure 10. Illustration of core components of CoPart by (park_copart_2019, 74).:
@@ 5.2.1. Proactive Approaches
Publications featuring approaches classified as proactive attempt to make resource management decisions in advance. They may enforce static partitioning schemes by profiling workloads and creating a fixed resource division or try to forecast future workload resource requirements based on past and current metrics.
@@
By profiling workloads as sensitive, or non-sensitive, to hardware prefetching of memory, (xiao_cppf_2019, 110) (xiao_cppf_2019, 110) propose a controller that aims to improve LLC management on modern systems equipped with hardware prefetching. They call their approach CPpf, which partitions LLC in accordance with the determined workload classifications. The system allocates a single exclusive cache way to each prefetching sensitive workload, leaving the rest to be shared among non-prefetching sensitive workloads.
@@
With the goal of predicting how a process’ instructions per cycle (IPC) metric changes as a result of a change in its available LLC, (kim_application_2019, 55) (kim_application_2019, 55) propose POCAT. POCAT is a trained machine learning model that can predict changes to IPC with an error of 4.7% without affecting current LLC allocations.
@@
To get a better understanding of how runtimes of BE high-performance computing (HPC) workloads are affected by core and LLC allocations, (aupy_co-scheduling_2019, 5) (aupy_co-scheduling_2019, 5) create a mathematical model that predicts how job execution times are affected by changes in scarce shared resource availability. Using their model, they are able to more efficiently partition resources among workloads to increase the efficiency of a system.
@@
A simple heuristic algorithm to partition LLC among VNFs is proposed by (durner_towards_2019, 26) (durner_towards_2019, 26). Their algorithm distributes available cache ways based on CPU utilization, which results in lowered total system utilization, increasing energy efficiency and QoS. However, only static configurations are considered where virtual network functions (VNFs) and their loads do not change. This makes their algorithm valid for very specific use cases only.
@@
In an attempt to make LLC partitioning more effective, Ginseng (funaro_ginseng_2016, 31) allow consumers (i.e. workloads) to bid on LLC space using Vickrey-Clarke-Groves auctioning. This allocates LLC to the process that would see the largest economic benefit from having said space. As the auctioning and allocation cycle is kept short, processes may acquire more LLC space at critical phases in its operation, while seceding it at times of less sensitivity.
@@
Public cloud providers aim to reduce their total cost of hardware ownership by optimizing its utilization. Making a VM scheduler LLC aware allows the provider to give LLC QoS guarantees to their customers as LLC ways may be exclusively, or partially exclusively, allocated to specific VMs. This is the goal of CacheSlicer, a VM scheduler proposed by (shahrad_provisioning_2021, 93) (shahrad_provisioning_2021, 93) that schedules VMs onto nodes based on low-cost, normal, and cache-sensitive VM categories.
@@
In public clouds, where different clients’ VMs share physical hosts, malicious actors may compromise security by extracting sensitive information from a co-located guest by manipulation of the LLC. CATalyst, as proposed by (liu_catalyst_2016, 60) (liu_catalyst_2016, 60), is a module that integrates into hypervisors that protect against such LLC side-channel attacks by partitioning the LLC to provide security guarantees.
@@
(zeng_playing_2021, 119) find that a simple static allocation of LLC is not sufficient to improve system performance when software prefetching of memory is used (zeng_playing_2021, 119). They show that improvements can be made even with a simple greedy partitioning scheme.
@@
An observation that allocating more LLC to an LC application may actually bring worse performance is made by (chung_enforcing_2019, 20) (chung_enforcing_2019, 20). This is due to the memory controller being overwhelmed by the requests of the remaining workloads that miss LLC more often due to less remaining usable after partitioning. The paper details their approach to mitigate this issue and the “memory virtual channels” they use to prioritize and order memory requests such that the LC application is guaranteed prioritized access to main memory as well as LLC.
@@
(thomas_dark_2018, 101) observe that as networks approach transfer speeds of 100 gigabits per second and beyond, the memory system becomes a bottleneck that can a cause large number of packet drops if incorrectly utilized (thomas_dark_2018, 101). As an attempt to mitigate this issue, they propose CacheBuilder. Cache builder is a system that, among other techniques, utilizes cache allocation to provide applications with limited control of the memory hierarchy such that packet drops may be minimized.
@@
By splitting LLC into two partitions, (sprabery_novel_2017, 96) (sprabery_novel_2017, 96) attempts to protect sensitive applications against cache side-channel attacks. Sensitive workloads are assigned to one, while other, insensitive, workloads share the other. The cores assigned to the sensitive partition will upon context switching perform a “state cleansing”. This means that the LLC used by the previous execution context will be flushed before the new context begins execution. This approach protects sensitive data at the cost of performance.